# Can We Detect Failures Without Failure Data? Uncertainty-Aware Runtime Failure Detection for Imitation Learning Policies

### v1

@@ VI-C Do failure detections align with human intuition?
Environment-dependent thresholds aid performance on robot hardware tasks. For simulation tasks (Fig. 4), ID-only bands computed once on successful ID rollouts work well for failure identification in both the ID and the OOD camera bump scenarios. This setting is practically preferable as it does not require collecting successful rollouts in each new environment to calibrate the failure detection threshold. However, on robot hardware tasks with OOD initial conditions (Figs. 5 and 11), most methods show degraded performance with ID-only bands, sometimes yielding zero TNR due to all rollouts being marked as failure. This over-conservative behavior likely occurs because OOD successful trajectories are less performant. We find on-robot policies to be more sensitive to the OOD environmental changes we employed than the simulation-based policies in response to the camera bump. Even when tasks are completed successfully in OOD conditions, trajectories often exhibit poor quality (e.g., slow execution, jitter) and higher (worse) scores compared to ID successful rollouts, suggesting they should potentially be classified as failures.
@@ VII Limitations
With FAIL-Detect, we demonstrate the promising potential of a failure detection method without access to failure data. The proposed two-staged method does have several limitations, which offer possible avenues for future work. We observed that, at times, the learned scores focus on simpler robot state information (e.g., gripper closed or open) over the higher-dimensional visual features. Finding score architectures that maximally make use of the visual information could improve failure detection performance. Additionally, our score candidates do not consider long sequences of temporal data. Although the CP band is temporally-aware, having the learned scores distill temporal patterns from historical time series data may lead to better failures predictions. Lastly, the timely detection of failures may be further improved by considering multimodal sensory information such as sound or tactile.
@@ VIII Conclusion
We present a modular two-stage uncertainty-aware runtime failure detection approach for generative imitation learning in manipulation tasks. Our method combines a learned scalar score and time-varying conformal prediction to accurately and efficiently identify failures while providing statistical guarantees. Through extensive experiments on diverse tasks, we show that the novel logpZO achieves the most consistently high performance among the considered candidates. Crucially, we demonstrate that unseen failures can be effectively detected without access to failure data, which may be expensive and difficult to collect to cover all possible failure scenarios. Our results highlight the potential of our method to enhance the safety and reliability of robotic systems in real-world applications.
@@ Acknowledgement
This work was primarily done during internship at Toyota Research Institute (TRI). We thank Vitor Guizilini, Benjamin Burchfiel, and David Snyder for the helpful discussions and feedback. We thank the robot teacher team at the TRI headquarters for the data collection, specifically Derick Seale and Donovon Jackson.

### v3

@@ VI-C Do failure detections align with human intuition?
Environment-dependent thresholds aid performance on robot hardware tasks. For simulation tasks (Fig. 4), ID-only bands computed on successful ID rollouts work well for failure identification in both the ID and the OOD camera bump scenarios. This setting is practically preferable as it does not require collecting successful rollouts in each new environment to calibrate the failure detection threshold. However, on-robot tasks with OOD initial conditions (Figs. 5, 13 and 14), most methods show degraded performance with ID-only bands, at times yielding low or close to zero TNR due to most rollouts being marked as failure. This over-conservative behavior likely occurs because OOD successful trajectories are less performant. We find on-robot policies to be more sensitive to the environmental changes we employed than the simulation-based policies in response to the camera bump. Even when tasks are completed successfully in these challenging conditions, trajectories often exhibit poor quality (e.g., slow execution, jitter) and higher (worse) scores compared to ID successful rollouts, suggesting they should potentially be classified as failures.
@@ VII Limitations
With FAIL-Detect, we demonstrate the promising potential of a failure detection method without access to failure data. The proposed two-staged method does have several limitations, which offer possible avenues for future work. We observed that, at times, the learned scores focus on simpler robot state information (e.g., gripper closed or open) over the higher-dimensional visual features. Finding score architectures that maximally make use of the visual information could improve failure detection performance. Meanwhile, we acknowledge the existence of false positives, especially in OOD settings where the trajectories degrade in performance. Our proposed setting-dependent CP band is meant to account for these environmental and behavioral shifts. However, more advanced techniques such as adaptively adjusting the CP significance level during inference could be helpful. Additionally, our score candidates do not consider long sequences of temporal data. Although the CP band is temporally-aware, having the learned scores distill temporal patterns from historical time series data may lead to better failures predictions. Lastly, the timely detection of failures may be further improved by considering multimodal sensory information such as sound or tactile.
@@ VIII Conclusion
We present a modular two-stage uncertainty-aware runtime failure detection approach for generative imitation learning in manipulation tasks. Our method combines a learned scalar score and time-varying conformal prediction to accurately and efficiently identify failures while providing statistical guarantees. Through extensive experiments on diverse tasks, we show that the novel logpZO score consistently achieves the highest performance among the considered candidates. Crucially, we demonstrate that unseen failures can be effectively detected without access to failure data, which may be difficult to collect to cover all possible failure scenarios. Our results highlight the potential of our method to enhance the safety and reliability of robotic systems in real-world applications.
@@ Acknowledgement
This work was primarily done during an internship at Toyota Research Institute (TRI). We thank Vitor Guizilini, Benjamin Burchfiel, and David Snyder for the helpful discussions and feedback. We thank the robot teacher team at the TRI headquarters for the data collection, specifically Derick Seale and Donovon Jackson.